### v2

@@ 5 Discussion
Another possible companion for a neutron star is a massive star that continuously emits stellar winds. As the neutron star interacts with the stellar wind of its companion, it can create activity windows that vary with frequency. To explain the observed DM of FRB 20121102A using these models, the FRB would need to be emitted when the neutron star is near the periastron of its orbit. This scenario would result in a variation of four orders of magnitude in the DM within approximately 10% of the orbital phase (Lyutikov et al., 2020). However, observations indicate that the activity window of FRB 20121102A spans about half of its cyclic period, and no DM variation of that magnitude has been observed.
@@ 6 Conclusions
We reported 7 observing epochs of FRB 20121102A monitoring and find no detections. We combined our observations with publicly available data of the source and divided it per frequency band and found a new best periodicity for the cyclic activity window to be days at L-band. This updated periodicity is more precise and with higher significance than the previously reported in Cruces et al. (2021b), mainly because of the larger and extended dataset. We also report the same periodicity using a normalised rates model that is more precise but has less significance than the one obtained by the binary model. We found a consistent periodicity at S-band and report that this active phase seems to end 19 days before the L-band active phase. No conclusions can be drawn from the C-band observations given the lack the observing epochs and in particular of detections. We strongly encourage the community to report both detections and non-detections of repeating FRBs and the full observation details. Sky and instrumentation statistics such as exposure time, observation duration, and telescope properties are fundamental to further constrain the nature of FRBs.

### v3

@@ 5 Discussion
Another possible companion for an NS is a massive star that continuously emits stellar winds. As the NS interacts with the stellar wind of its companion, it can create activity windows that vary with frequency. To explain the observed DM of FRB 20121102A using these models, the FRB would need to be emitted when the NS is near the periastron of its orbit. This scenario would result in a variation of four orders of magnitude in the DM within approximately 10% of the orbital phase (Lyutikov et al., 2020). However, observations indicate that the activity window of FRB 20121102A spans about half of its cyclic period, and no DM variation of that magnitude has been observed.
@@ 6 Conclusions
We report seven observing epochs of FRB 20121102A monitoring and find no detections. We combined our observations with publicly available data on the source, dividing this combined dataset by frequency band, and find a new best periodicity for the cyclic activity window to be days at L-band. This updated periodicity is more precise and has a higher significance than that previously reported in Cruces et al. (2021b), mainly because of the larger and extended dataset. We also report the same periodicity using a normalised rates model that is more precise but has a lower significance than the one obtained by the binary model. We find a consistent periodicity in the S band and report that this active phase seems to end 19 days before the L-band active phase. No conclusions can be drawn from the C-band observations given the lack of observing epochs and, in particular, of detections. We strongly encourage the community to report both detections and non-detections of repeating FRBs, along with full observation details. Sky and instrumentation statistics such as exposure time, observation duration, and telescope properties are fundamental to further constraining the nature of FRBs.